# GeSubNet: Gene Interaction Inference for Disease Subtype Network Generation

### v1

@@ 4.2 Experiment-I: network inference
Objective: This experiment evaluates the effectiveness of subtype-specific networks, following our Hypothesis-1: (1) $|\mathcal{G}|_{y}\ll|\mathcal{G}|$, ensuring the generated network is sparse compared to the original; (2) each subtype network exhibits structural differences from the others.
Setup and Metrics: We train GenSubNet for each cancer (the parameter settings can be found in Appendix D), and then evaluate the generated graphs for subtypes on two factors:
  • Sparsity Assessment: we utilize the Coefficient of Degree Variation (CDV) (Pržulj, 2007) to measure the variability in gene nodes within a network. A higher CDV value indicates that most genes have very few interactions (edges). Thus, GenSubNet infers that the network becomes sparser because only a few active genes dominate the interactions in this subtype network.
@@
Results. Table 2 presents GenSubNet significantly outperforms all baseline methods in terms of GED, DCS, and CDV metrics across four cancer types. Compared with the second-best baseline, CSGNN, GenSubNet achieves improvements of 35.8%/32.4%/20.2%/34.1% in terms of GED across all four tasks. Additionally, it delivers a relative reduction of 29.8%/13.5%/21.6%/19.3% in terms of DCS. For CDV, the improvements are 33.4%/13.7%/17.9%/15.3%, respectively. In summary, when evaluating BRCA, GBM, LGG, and OV, GenSubNet consistently achieves lower DCS scores and higher GED and CDV scores. This indicates that the generated subtype-specific gene networks are sparse but structurally unique, i.e., they are significantly different from each other. The OV results are apparently unsatisfactory, but this aligns with existing knowledge (Lawler et al., 2017) that OV is a challenging cancer type due to the limited available samples (only 291 patients in Table 1) and the lack of information on their pathogenic mechanisms in existing knowledge databases.
Table 2: Baseline comparison results on GED, DCS, and CDV for the proposed and all compared methods. GED, DCS, and CDV are subjected to a min-max normalization. The best-performing results are highlighted in bold. The second-best results are highlighted in underline.
@@
WGCNA 0.42 0.39 0.83 0.43 0.47 0.83 0.45 0.53 0.82 0.24 0.25 0.83
wTO 0.44 0.43 0.79 0.45 0.47 0.83 0.43 0.59 0.76 0.26 0.25 0.83
ARACNe 0.47 0.45 0.73 0.44 0.43 0.79 0.43 0.57 0.76 0.23 0.25 0.81
LEAP 0.49 0.44 0.78 0.48 0.45 0.78 0.44 0.55 0.77 0.22 0.24 0.84
GAERF 0.54 0.58 0.64 0.46 0.48 0.76 0.55 0.56 0.83 0.34 0.36 0.82
LR-GNN 0.54 0.59 0.62 0.57 0.61 0.75 0.56 0.66 0.72 0.34 0.37 0.82
CSGNN 0.65 0.66 0.52 0.65 0.64 0.74 0.58 0.68 0.73 0.35 0.35 0.80
GenSubNet 0.75 0.78 0.47 0.73 0.74 0.67 0.67 0.74 0.62 0.45 0.44 0.75
@@ 4.3 Experiment-II: Biological Meaningfulness
Objective: While three graph metrics show the statistical significance of the generated network, this experiment further evaluates their biological relevance, following our Expectation-2. (1) Instead of structural differences, we further assess whether each network shows biologically functional differences from other networks. (2) We examine whether the generated networks have the potential to narrow down key genes that contribute more specifically to their respective subtypes.
Setup and Metrics: We hence conduct two experiments as follows:
  • Gene Ontology (GO) Analyses (Ashburner et al., 2000): This method counts the number of unique GO terms associated with the genes in each network. GO terms describe gene functions across biological processes, molecular functions, and cellular components, enabling comparisons between gene networks. For example, if $GO(\mathcal{G}_{1}):=\{A,B,C\}$ and $GO(\mathcal{G}_{2}):=\{A,D,E\}$, where $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ represent two generated subtype gene networks. $GO(\cdot)$ denotes the sets of GO terms for two networks. Here the number of Enriched Biological Functions (#EBF) is 4, i.e., $\{B,C,D,E\}$, since $A$ is the shared GO term. We evaluate GO for each cancer dataset across all baselines. A high #EBF value indicates greater functional diversity and biological differences between subtypes.
  • Simulated Gene Knockout: This is a computational technique that mimics the effects of gene knockout experiments without physically altering the genome (Bergman & Siegal, 2003). In this simulation, a gene is either deleted or deactivated to study its role within a specific subtype by observing changes in the patient sample distribution. As we described in an observation in Sec. 2, the key genes with significant expression differences form a small, limited set (Goh et al., 2007a), which leads to a distribution shift in patient samples during simulation experiments.
@@
    where $T$ is the total number of knockout tests, $n$ is the number of patient samples within a subtype, $x_{i}$ represents an individual patient sample, $k$ is a scaling factor (e.g., 1.0 or 1.5) used to adjust the threshold, and $\sigma_{j}$ is the standard deviation of sample distances. Notably, this metric is only used after model training and cannot involved in modeling training. More details on the simulated Gene Knockout can be found in Appendix G.
Figure 3: The Venn diagrams illustrate the overlap in GO terms resulting from different methods (WGCNA, CSGNN, LR-GNN, and GenSubNet) across four cancers. Shared and unique function items are listed here. A full list is provided in Appendix F. We highlight some unique function items that are well-supported by biological evidence in bold.
Results. Table 3 presents the GO analysis results, where our method consistently achieves the highest number of enriched biological functions (#EBF) across all datasets. These higher values indicate that the generated networks not only exhibit structural differences but also show functional distinctions from others from a biological perspective.
Figure 3 presents Venn diagrams of detailed GO analysis for four cancer datasets, highlighting overlaps/unique in biological functions among three selected baselines and our method. GenSubNet consistently identifies several unique functions across all datasets, while other methods rarely uncover unique functions, even when they achieve a comparable #EBF. For instance, in the LGG cancer dataset, CSGNN identifies 4 #EBF but finds no unique functions, whereas GenSubNet identifies 6 #EBF with 3 unique functions. From a biological perspective, GenSubNet demonstrates a robust array of enriched GO terms across different cancers, including pathways like Apoptotic signaling, Wnt signaling, Tumor necrosis factor signaling, and Cell proliferation. These terms represent critical cancer-related biological functions common to many cancers (Aktipis & Nesse, 2013), as shown in Table 8 in Appendix F. For unique functions, GenSubNet identified the ”Immune diseases” function in BRCA, which has evident support as being related to breast cancer (McAlpine et al., 2012), and the ”DNA damage checkpoint signaling” pathway, which is specific to GBM (Cheng et al., 2011).
Table 3: The comparison results on #EBF between GenSubNet and the baselines. Only biological functions with high statistical significance (p-value $<0.05$) are reported.
@@
GenSubNet 8 6 6 5
Figure 4 illustrates the results of the simulated gene knockout experiments. Subfigure (a) visualizes an example of patient distribution before (red-marked points) and after (green-marked points) the Simulated Gene Knockout in both target and control groups (subtypes). In the left subfigure, there are almost no differences between the before and after distributions for the low-ranking gene set. In contrast, the right subfigure shows a significant shift in patient distribution, indicating that the suppression of high-ranking genes has a greater impact.
Figure 4(b) provides a statistical summary of the results across 11,327 genes in BRCA. GenSubNet achieves the highest $\Delta_{\text{SR}}$ for high-ranking genes, with an 83% likelihood of significantly shifting sample distributions. Meanwhile, the 12% $\Delta_{\text{SR}}$ for low-ranking genes suggests that GenSubNet effectively filters out common genes. Other methods show much lower $\Delta_{\text{SR}}$ values for high-ranking genes, ranging from 20%-30%, nearly matching those for low-ranking genes. Notably, while GNN-based methods like LR-GNN and CSGNN achieve comparable results in graph statistical metrics, their biological relevance is lower. This discrepancy arises because their objective functions aim only to reconstruct general disease networks, including irrelevant gene interactions, for all subtype samples. Although gene expression data embeddings result in different graph structures, these methods do not explicitly learn the distinct gene interactions unique to disease subtypes. However, learning a representation that incorporates prior knowledge while explicitly distinguishing patient samples in different subtypes is the key focus of this paper. This simulation experiment further validates the effectiveness of our method and demonstrates that GenSubNet maintains biological significance.
@@
Figure 5: The obtained gene networks for two BRCA patient groups: the Normal-like group (network A) and the Basal-like group (network B). Comparisons were made between our method and two baseline methods, CSGNN and WGCNA.
Figure 6: (1) expression level distributions and (2) the expression heatmaps of the top-3 genes from the high-ranking and low-ranking gene sets among different patient groups. Different colors in heat maps indicate the gene expression level.
@@ 4.4 Case Study
The case study on BRCA follows established protocols in bioinformatics gene function studies (Huang et al., 2009). The analysis workflow is available in Appendix  H.1 and  H.2.
Figure 5 shows the gene networks A and B obtained for two BRCA subtypes. We observe that GenSubNet generated gene networks with more distinct gene nodes. The networks show significant differences between the two subtypes, whereas the baselines produce more similar networks.
Figure 6(1) presents the gene expression distribution for the high-ranking and low-ranking gene sets. Different patient groups are marked in various colors to represent the ground truth. In the first column, we observe minimal differences in the expression distribution of low-ranking genes across patient groups. However, significant differences are evident in the high-ranking gene sets, as shown by the noticeable shift in distribution peaks.
@@ 5 Conclusions
This paper introduced GenSubNet, a framework for inferring disease subtype-specific gene networks. GenSubNet includes sample and gene embedding learning modules that capture the characteristics of both patients and the prior gene graph. These embeddings are then utilized to reconstruct the input gene profile in the network inference module. This approach incorporates patient group information into the updated gene embeddings, enabling more accurate gene network inference specific to patient groups. As a result, GenSubNet offers a unified framework for group-specific gene network inference on real-world clinical data. Importantly, we demonstrated the reliability of GenSubNet through a series of biological validations. We believe that GenSubNet will be a valuable tool for disease research and other gene function-related applications.
@@ B.2 Preprocessing of gene expression data
data <- read.csv("gene_expression_data.csv")
@@ Appendix C Baselines
(1) Weighted Gene Co-expression Network Analysis (WGCNA) (Langfelder & Horvath, 2008) utilizes Pearson correlation to identify modules of highly correlated genes, where genes within the same module are likely to be functionally related or involved in similar biological processes. (2) Weighted Topological Overlap (wTO) (Gysi et al., 2018) normalizes the chosen correlation by all other correlations and calculates a probability for each edge in the network. (3) Algorithm for the Reconstruction of Accurate Cellular Networks (ARACNe) (Margolin et al., 2006) calculates the mutual information between pairs of nodes and then removes indirect relationships during network building. (4) Lag-based Expression Association for Pseudotime-series (LEAP) (Specht & Li, 2017) utilizes pseudotime ordering to infer the directionality between genes in the network. (5) Graph Auto-encoder and Random Forest (GAERF) (Wu et al., 2021) learns features of nodes by a graph auto-encoder and concatenates features of two nodes as input for the random forest classifier. (6) Link Representation-Graph Neural Network (LR-GNN) (Kang et al., 2022) generates embeddings using a GCN encoder and then applies a propagation rule to create link representations for predicting associations in networks. (7) Contrastive Self-supervised Graph Neural Network (CSGNN) (Zhao et al., 2021) predicts node interactions in networks by employing a mix-hop aggregator and a contrastive self-supervised GNN. WGCNA, wTO, ARACNe, and LEAP are well-used traditional methods that use only non-graph gene expression data as input, while GAERF, LR-GNN, and CSGNN are deep learning-based methods that use known paired integrations or networks as input. These methods are reported to have competitive performance for similar tasks like molecular interaction prediction. It is also worth noting that these methods tend to perform better when supplementary data, such as sequence data, is available.

### v3

@@ 4.2 Experiment-I: Network Inference
Objective. This experiment evaluates the effectiveness of subtype-specific networks, following our Hypothesis-1: (1) $|\mathcal{G}|_{y}\ll|\mathcal{G}|$, ensuring the generated network is sparse compared to the original; (2) each subtype network exhibits structural differences from the others.
Setup and Metrics. We train GeSubNet for each cancer (the parameter settings can be found in Appendix D), and then evaluate the generated graphs for subtypes on two factors:
  • Sparsity Assessment: we utilize the Coefficient of Degree Variation (CDV) (Pržulj, 2007) to measure the variability in gene nodes within a network. A higher CDV value indicates that most genes have very few interactions (edges). Thus, GeSubNet infers that the network becomes sparser because only a few active genes dominate the interactions in this subtype network.
@@
Results. Table 2 presents GeSubNet significantly outperforms all baseline methods in terms of GED, DCS, and CDV metrics across four cancer types. Compared with the second-best baseline, CSGNN, GeSubNet achieves improvements of 35.8%/32.4%/20.2%/34.1% in terms of GED across all four tasks. Additionally, it delivers a relative reduction of 29.8%/13.5%/21.6%/19.3% in terms of DCS. For CDV, the improvements are 33.4%/13.7%/17.9%/15.3%, respectively. In summary, when evaluating BRCA, GBM, LGG, and OV, GeSubNet consistently achieves lower DCS scores and higher GED and CDV scores. This indicates that the generated subtype-specific gene networks are sparse but structurally unique, i.e., they are significantly different from each other. The OV results are apparently unsatisfactory, but this aligns with existing knowledge (Lawler et al., 2017) that OV is a challenging cancer type due to the limited available samples (only 291 patients in Table 1) and the lack of information on their pathogenic mechanisms in existing knowledge databases.
Table 2: Baseline comparison results on GED, DCS, and CDV for the proposed and baselines. GED, DCS, and CDV are subjected to min-max normalization. The best-performing results are highlighted in bold. The second-best results are highlighted in underline.
@@
WGCNA 0.42 ± 0.02 0.39 ± 0.03 0.83 ± 0.04 0.43 ± 0.02 0.47 ± 0.03 0.83 ± 0.04 0.45 ± 0.03 0.53 ± 0.02 0.82 ± 0.04 0.24 ± 0.02 0.25 ± 0.03 0.83 ± 0.04
wTO 0.44 ± 0.02 0.43 ± 0.02 0.79 ± 0.03 0.45 ± 0.02 0.47 ± 0.02 0.83 ± 0.04 0.43 ± 0.03 0.59 ± 0.03 0.76 ± 0.04 0.26 ± 0.02 0.25 ± 0.03 0.83 ± 0.04
ARACNe 0.47 ± 0.02 0.45 ± 0.03 0.73 ± 0.03 0.44 ± 0.02 0.43 ± 0.02 0.79 ± 0.03 0.43 ± 0.03 0.57 ± 0.03 0.76 ± 0.04 0.23 ± 0.02 0.25 ± 0.03 0.81 ± 0.03
LEAP 0.49 ± 0.03 0.44 ± 0.03 0.78 ± 0.03 0.48 ± 0.03 0.45 ± 0.03 0.78 ± 0.03 0.44 ± 0.03 0.55 ± 0.02 0.77 ± 0.04 0.22 ± 0.02 0.24 ± 0.03 0.84 ± 0.04
GAERF 0.54 ± 0.06 0.58 ± 0.07 0.64 ± 0.05 0.46 ± 0.04 0.48 ± 0.06 0.76 ± 0.05 0.55 ± 0.05 0.56 ± 0.06 0.83 ± 0.07 0.34 ± 0.05 0.36 ± 0.06 0.82 ± 0.06
LR-GNN 0.54 ± 0.05 0.59 ± 0.06 0.62 ± 0.04 0.57 ± 0.06 0.61 ± 0.07 0.75 ± 0.05 0.56 ± 0.06 0.66 ± 0.07 0.72 ± 0.05 0.34 ± 0.05 0.37 ± 0.06 0.82 ± 0.07
CSGNN 0.65 ± 0.06 0.66 ± 0.07 0.52 ± 0.06 0.65 ± 0.07 0.64 ± 0.06 0.74 ± 0.05 0.58 ± 0.06 0.68 ± 0.07 0.73 ± 0.06 0.35 ± 0.05 0.35 ± 0.06 0.80 ± 0.05
GeSubNet 0.75 ± 0.04 0.78 ± 0.04 0.47 ± 0.05 0.73 ± 0.04 0.74 ± 0.05 0.67 ± 0.04 0.67 ± 0.05 0.74 ± 0.04 0.62 ± 0.05 0.45 ± 0.04 0.44 ± 0.04 0.75 ± 0.04
@@ 4.3 Experiment-II: Biological Meaningfulness
Objective. While three graph metrics show the statistical significance of the generated network, this experiment further evaluates their biological relevance, following our Hypothesis-2. (1) Instead of structural differences, we further assess whether each network shows biologically functional differences from other networks. (2) We examine whether the generated networks have the potential to narrow down key genes that contribute more specifically to their respective subtypes.
Setup and Metrics. We hence conduct two experiments as follows:
  • Gene Ontology (GO) Analyses (Ashburner et al., 2000a): This method counts the number of unique GO terms (under the category Biological Process (Desmedt et al., 2008)) associated with the genes in each network. GO terms describe gene functions across biological processes, molecular functions, and cellular components, enabling comparisons between gene networks. For example, if $GO(\mathcal{G}_{1}):=\{A,B,C\}$ and $GO(\mathcal{G}_{2}):=\{A,D,E\}$, where $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ represent two generated subtype gene networks. $GO(\cdot)$ denotes the sets of GO terms for two networks. Here the number of Enriched Biological Functions (#EBF) is 4, i.e., $\{B,C,D,E\}$, since $A$ is the shared GO term. We evaluate GO for each cancer dataset across all baselines. A high #EBF value indicates greater functional diversity and biological differences between subtypes.
  • Simulated Gene Knockout (Bergman & Siegal, 2003): This is a computational technique that mimics the effects of gene knockout experiments without physically altering the genome. In this simulation, a gene is either deleted or deactivated to study its role within a specific subtype by observing changes in the patient sample distribution. As we described in an observation in Sec. 2, the key genes with significant expression differences form a small, limited set (Goh et al., 2007a), which leads to a distribution shift in patient samples during simulation experiments.
@@
    where $T$ is the total number of knockout tests, $n$ is the number of patient samples within a subtype, $x_{i}$ represents an individual patient sample, $k$ is a scaling factor (e.g., 1.0 or 1.5) used to adjust the threshold, and $\sigma_{j}$ is the standard deviation of sample distances. Notably, this metric is only used after model training and cannot be involved in modeling training. More details on the simulated Gene Knockout can be found in Appendix H.
Figure 3: The Venn diagrams illustrate the overlap in GO terms resulting from different methods (WGCNA, CSGNN, LR-GNN, and GeSubNet) across four cancers. Shared and unique function items are listed here. A full list is provided in Appendix G. We highlight some unique function items that are well-supported by biological evidence in bold.
Results. Table 3 presents the results of the GO analysis , where our method consistently achieves the highest #EBF value across all datasets. These higher values indicate that the generated networks not only exhibit structural differences but also show functional distinctions from others from a biological perspective  (Ashburner et al., 2000b; Wu et al., 2021b).
Table 3: The comparison results on #EBF between GeSubNet and the baselines. Only biological functions with high statistical significance (p-value $<0.05$) are reported.
@@
GeSubNet 8 6 6 5
Figure 3 presents Venn diagrams of detailed GO analysis for four cancer datasets, highlighting overlaps and unique in biological functions among three selected baselines and our method. We can consistently identify several unique functions across all datasets, while other methods rarely uncover unique functions, even when they achieve a comparable #EBF. For instance, in the LGG cancer dataset, CSGNN identifies 4 #EBF but finds no unique functions, whereas our method identifies 6 #EBF with 3 unique functions. From a biological perspective, our method demonstrates a robust array of enriched GO terms across different cancers, including pathways like Apoptotic signaling, Wnt signaling, Tumor necrosis factor signaling, and Cell proliferation. These terms represent critical cancer-related biological functions common to many cancers (Aktipis & Nesse, 2013), as shown in Table 9 in Appendix G. For unique functions, our method identifies the “Immune diseases” function in BRCA, which has evident support as being related to breast cancer (McAlpine et al., 2012), and the “DNA damage checkpoint signaling” pathway, which is specific to GBM (Cheng et al., 2011).
Figure 4 shows the results of the simulated gene knockout experiments. The subfigure (a) visualizes an example of patient distribution before (red-marked points) and after (green-marked points) the Simulated Gene Knockout in both target and control groups (subtypes). In the left subfigure, there are almost no differences between the before and after distributions for the low-ranking gene set. In contrast, the right subfigure shows a significant shift in patient distribution, indicating that the suppression of high-ranking genes has a greater impact.
Figure 4(b) further provides a quantitative table across 11,327 genes in BRCA. our method achieves the highest $\Delta_{\text{SR}}$ for high-ranking genes, with an 83% likelihood of significantly shifting sample distributions. Meanwhile, the 12% $\Delta_{\text{SR}}$ for low-ranking genes suggests that our method effectively filters out common genes. Other baselines exhibit substantially lower $\Delta_{\text{SR}}$ values for high-ranking genes, ranging from 20%-30%, nearly matching those for low-ranking genes. Notably, while GNN-based methods like LR-GNN and CSGNN achieve comparable results in graph statistical metrics, their biological relevance is lower. This discrepancy arises because their objective functions aim only to reconstruct general disease networks, including irrelevant gene interactions, for all subtype samples. Although gene expression data embeddings result in different graph structures, these methods do not explicitly learn the distinct gene interactions unique to disease subtypes. However, learning a representation that incorporates prior knowledge while explicitly distinguishing patient samples in different subtypes is the key focus of this paper. This simulation experiment further validates the effectiveness of our method and demonstrates that GeSubNet maintains biological significance.
@@
Figure 5: The obtained gene networks for two BRCA patient groups: the Normal-like group (network A) and the Basal-like group (network B).
@@ 4.4 Case Study
The case study on BRCA cancer follows established protocols in bioinformatics gene function studies (Huang et al., 2009). The analysis workflow is available in Appendix  I.1 and  I.2.
Figure 5 shows the gene networks A and B obtained for two BRCA subtypes. We observe that our method can generate gene networks with more distinct gene nodes. The networks show significant differences between the two subtypes, whereas the baselines produce more similar networks.
Figure 6(1) shows the gene expression distribution for the high-ranking and low-ranking gene sets. Different patient groups are marked in various colors to represent the ground truth. In the first column, we observe minimal differences in the expression distribution of low-ranking genes across patient groups. However, significant differences are evident in the high-ranking gene sets, as shown by the noticeable shift in distribution peaks.
@@
Figure 6: (1) expression level distributions and (2) the expression heatmaps of the top-3 genes from the high-ranking and low-ranking gene sets among different patient groups.
@@ 5 Conclusions
This paper introduced GeSubNet, a framework for inferring disease subtype-specific gene networks. GeSubNet includes sample and gene embedding learning modules that capture the characteristics of both patients and the prior gene graph. These embeddings are then utilized to reconstruct the input gene profile in the network inference module. This approach incorporates patient group information into the updated gene embeddings, enabling more accurate gene network inference specific to patient groups. In general, we explored a method for group-specific gene network inference on real-world clinical data. Importantly, we demonstrated the reliability of GeSubNet through a series of biological validations. We believe that continued investigation will bridge computational and biological sciences, advancing the understanding of diseases and foundational gene roles.
6 Acknowledgment
This work was supported by JSPS KAKENHI Grant-in-Aid for Scientific Research Number JP24K20778, JST CREST JPMJCR23M3, NSF award SCH-2205289, SCH-2014438, IIS-2034479, and AIST policy-based budget projects of “Research DX Platforms”. This paper is based on results obtained from the project, “Research and Development Project of the Enhanced infrastructures for Post 5G Information and Communication Systems” (JPNP20017), commissioned by the New Energy and Industrial Technology Development Organization (NEDO).
@@ B.2 Preprocessing of Gene Expression data
data <- read.csv(‘‘gene_expression_data.csv")
@@ Appendix C Baselines
C.1 The Principle of Selecting Baseline Methods
We collected baselines from both statistical methods and GNN-based methods. For statistical methods that do not involve neural networks, we first included two of the most classical and widely used methods in the bio-network computing field: WGCNA (Langfelder & Horvath, 2008) and ARACNe (Margolin et al., 2006). These methods have continually contributed to bio-network studies for many years.
In addition, we want to include more recently proposed methods in this category. Our selection principles were as follows: (1) the method was recently published (and approximately a decade after the above classical methods), (2) it provides open-source code (available on platforms such as GitHub), (3) it is easy to use (with outputs compatible with popular analysis packages), and (4) it has been used as a baseline in related bio-network research. Based on these criteria, we selected two methods, wTO (Gysi et al., 2018) and LEAP (Specht & Li, 2017).
For the GNN-based methods, our principles aligned with the above five lines for selecting newly proposed statistical methods. We selected three methods (i.e., GAERF (Wu et al., 2021a), LR-GNN (Kang et al., 2022), and CSGNN (Zhao et al., 2021)) that meet these principles.
C.2 Baseline Methods
(1) Weighted Gene Co-expression Network Analysis (WGCNA) (Langfelder & Horvath, 2008) utilizes Pearson correlation to identify modules of highly correlated genes, where genes within the same module are likely to be functionally related or involved in similar biological processes. (2) Weighted Topological Overlap (wTO) (Gysi et al., 2018) normalizes the chosen correlation by all other correlations and calculates a probability for each edge in the network. (3) Algorithm for the Reconstruction of Accurate Cellular Networks (ARACNe) (Margolin et al., 2006) calculates the mutual information between pairs of nodes and then removes indirect relationships during network building. (4) Lag-based Expression Association for Pseudotime-series (LEAP) (Specht & Li, 2017) utilizes pseudotime ordering to infer the directionality between genes in the network. (5) Graph Auto-encoder and Random Forest (GAERF) (Wu et al., 2021a) learns features of nodes by a graph auto-encoder and concatenates features of two nodes as input for the random forest classifier. (6) Link Representation-Graph Neural Network (LR-GNN) (Kang et al., 2022) generates embeddings using a GCN encoder and then applies a propagation rule to create link representations for predicting associations in networks. (7) Contrastive Self-supervised Graph Neural Network (CSGNN) (Zhao et al., 2021) predicts node interactions in networks by employing a mix-hop aggregator and a contrastive self-supervised GNN. WGCNA, wTO, ARACNe, and LEAP are well-used traditional methods that use only non-graph gene expression data as input, while GAERF, LR-GNN, and CSGNN are deep learning-based methods that use known paired integrations or networks as input. These methods are reported to have competitive performance for similar tasks like molecular interaction prediction. It is also worth noting that these methods tend to perform better when supplementary data, such as sequence data, is available.